OpenStudy (anonymous):

PLEASE HELP!! what is the simplest form of y^2 - y - 12/ y^2 + 8y +15.

OpenStudy (anonymous):

1. y - 12 / 9y \[\times\] y +15

OpenStudy (anonymous):

here are the choices. y-4/y-5 y-5/y+5 y+4/y-5 y+4/y+5

OpenStudy (anonymous):

I think it's the 3rd one.

jhonyy9 (jhonyy9):

this will be (y-4)(y+3) ---------- = (y+5)(y+3) can you continue ,simplifie it ?

OpenStudy (anonymous):

i use FOIL right?

jhonyy9 (jhonyy9):

jackie -- check your choices !!! there not is the right answer sure

OpenStudy (anonymous):

yes i'm sure those are the answers.

jhonyy9 (jhonyy9):

the last will be right if you corrcte the --- oh not i see it on second you have wrote 5 where you need writing 4 in numerator so than will be right exactly what i have got to right answer

jhonyy9 (jhonyy9):

y-5 in numerator not can being never

jhonyy9 (jhonyy9):

there is (y-4)/(y-5)

jhonyy9 (jhonyy9):

good luck bye